6-16. The town of Hill Valley is issuing a 30-year bond with a face value of $50 million and a stated annual interest rate of 4 percent. The town will make interest payments twice a year. Calculate the semiannual interest payment. Calculate how much Hill Valley will receive from the bond offering under the following conditions: Market interest rates remain unchanged at the time of the offering. Market interest rates increase to 4.2 percent at the time of the offering
Answer #1

1: semiannual interest payment.= 4%*$50m/2 = $ 1 million

2: Price if the Market interest rate is 4% = $50 million

(Bond will be sold at par if the interest rate equals coupon rate)

3: Price if the Market interest rate is 4.2% = $48.3 million

Using financial calculator
Input: FV = 50, PMT=1, I/Y=4.2/2=2.1 , N=30*2=60

Solve for PV as -48.3